# Normal Quaternionic Matrices and Finitely Generated Witt Rings

### v1

@@ Appendix A Lists of Normal Quaternionic Matrices
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{3}\Delta$ (0,0,0,0,0,1,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,0,1,0,0,0,2,0,4,0,8,0,0) $\mathbb{L}_{1,0}\Delta_{2}\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,0,1,0,0,0,2,4,8,0,16,0,0) $\mathbb{L}_{6,1}$ (0,0,0,0,0,1,0,0,1,0,1,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{4,1}$ (0,0,0,0,0,1,0,0,1,0,2,0,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{3}\Delta$ (0,0,0,0,0,1,0,0,2,0,0,4,0,8,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{4,0}$ (0,0,0,0,0,1,0,0,2,0,2,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta^{2}\times\mathbb{L}_{1,1}\Delta$ (0,0,0,0,0,1,0,0,2,0,4,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,0,1,0,0,2,0,4,0,8,0,16,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,0,1,0,0,2,0,4,8,0,16,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{2}\Delta_{2}$ (0,0,0,0,0,1,0,2,4,0,8,16,0,32,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}\Delta_{3}$ (0,0,0,0,0,1,2,4,8,0,16,32,0,64,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{3}\Delta$ (0,0,0,0,1,2,0,0,0,2,0,0,4,0,8,0) $\mathbb{L}_{1,1}^{2}\Delta^{2}$ (0,0,0,0,1,2,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{4,1}$ (0,0,0,0,1,2,0,0,0,2,1,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta^{2}$ (0,0,0,0,1,2,0,0,0,2,4,0,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,1,2,0,0,0,2,4,0,8,0,16,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,0,0,1,2,0,0,0,4,8,0,0,0,0,16) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{2}\Delta_{2}$ (0,0,0,0,1,2,0,0,4,8,16,0,0,0,0,32) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,0,1,2,0,4,0,0,8,0,0,0,0,16) $\mathbb{L}_{1,0}\Delta_{2}\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,0,1,2,4,8,0,0,16,0,0,0,0,32) $\mathbb{L}_{1,1}^{2}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,1,2,4,0,0,0,4,0,0,8,16,0,0) $\mathbb{L}_{1,1}\Delta^{3}$ (0,0,0,1,2,4,0,0,0,4,0,2,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,0,1,2,4,0,0,0,4,0,8,0,16,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{2}\Delta_{2}$ (0,0,0,1,2,4,0,0,0,4,8,16,0,32,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta_{3}$ (0,0,0,1,2,4,8,0,0,0,0,0,16,32,64) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}\Delta_{3}$ (0,0,1,2,4,8,0,0,0,8,16,32,0,64,0,0) $\mathbb{L}_{1,1}\Delta_{2}^{2}$ (0,0,1,2,4,8,0,0,4,8,16,0,0,0,0,32)
normal quaternionic matrices of order 6, level $>2$
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{3}\times\mathbb{L}_{1,0}^{2}\Delta$ (1,0,0,0,0,0,0,0,0,1,0,2,0,4,0,0) $\mathbb{L}_{3}\times\mathbb{L}_{1,0}\Delta_{2}$ (1,0,0,0,0,0,0,0,0,1,2,4,0,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (1,0,0,0,0,0,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (1,0,0,0,0,0,0,0,0,2,4,8,0,16,0,0) $\mathbb{L}_{3}\times\mathbb{L}_{1,1}^{2}\Delta$ (1,0,0,0,0,2,0,0,0,2,0,0,4,1,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (1,0,0,0,0,2,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (1,0,0,0,0,2,0,0,0,2,0,4,0,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (1,0,0,0,0,2,0,0,0,2,4,8,0,16,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,0,2,0,0,0,4,0,0,8,16,0,0) $\mathbb{L}_{1,0}^{2}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,0,2,0,0,0,4,0,8,0,16,0,0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta\times\mathbb{L}_{1,0}\Delta^{2}$ (1,0,0,0,0,2,0,0,0,4,8,16,0,32,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{3}\Delta$ (1,0,0,0,0,2,0,0,1,4,8,0,0,0,0,16) $\mathbb{L}_{1}\times\mathbb{L}_{5}$ (1,0,0,0,0,2,0,0,2,0,2,0,0,0,0,0) $\mathbb{L}_{3}^{2}$ (1,0,0,0,0,2,0,0,2,0,3,0,0,0,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{3}\times\mathbb{L}_{1,0}\Delta$ (1,0,0,0,0,2,0,0,2,0,4,0,0,0,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{4,0}$ (1,0,0,0,0,2,0,0,4,0,4,0,0,0,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,0}\Delta^{2}$ (1,0,0,0,0,2,0,0,4,0,8,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (1,0,0,0,0,2,0,0,4,0,8,0,16,0,32,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (1,0,0,0,2,4,0,0,0,4,0,0,8,2,0,0) $\mathbb{L}_{1,1}^{2}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,4,0,0,0,4,0,0,8,16,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,4,0,0,0,4,0,8,0,16,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{3}\Delta$ (1,0,0,0,2,4,0,0,0,4,1,8,0,16,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{3}$ (1,0,0,0,2,4,0,0,0,4,2,0,0,0,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{4,1}$ (1,0,0,0,2,4,0,0,0,4,4,0,0,0,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta$ (1,0,0,0,2,4,0,0,0,4,8,0,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (1,0,0,0,2,4,0,0,0,4,8,16,0,32,0,0) $\mathbb{L}_{3}\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,4,0,0,0,8,2,0,0,0,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,4,0,0,0,8,16,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2})\Delta$ (1,0,0,0,2,4,0,0,0,8,16,0,0,0,0,32) $\mathbb{L}_{3}\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,0,0,2,4,0,0,2,4,1,0,0,0,0,8) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,0,0,2,4,0,0,2,4,8,0,0,0,0,16) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (1,0,0,0,2,4,0,0,2,8,16,0,0,0,0,32) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{2}$ (1,0,0,0,2,4,0,0,8,16,32,0,0,0,0,64) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,1}\Delta^{2}$ (1,0,0,2,4,8,0,0,0,8,0,4,0,0,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,2,4,8,0,0,0,8,0,16,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2})\Delta$ (1,0,0,2,4,8,0,0,0,8,0,16,0,32,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,0,2,4,8,0,0,0,8,2,4,0,16,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (1,0,0,2,4,8,0,0,0,8,2,16,0,32,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{2}$ (1,0,0,2,4,8,0,0,0,8,16,32,0,64,0,0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta^{2}$ (1,0,0,2,4,8,0,0,0,16,0,32,0,0,8,0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,0,2,4,8,0,0,0,16,2,4,0,32,0,0) $\mathbb{L}_{1}^{3}\times\mathbb{L}_{3}$ (1,0,0,2,4,8,2,0,0,0,0,0,0,0,0,0) $\mathbb{L}_{3}\times\mathbb{L}_{1}^{2}\Delta$ (1,0,0,2,4,8,2,0,0,0,0,0,0,0,0,16)
[…]

### v2

@@ Appendix A Lists of Normal Quaternionic Matrices
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{3}\Delta$ (0,0,0,0,1; 0,0,0,2; 0,0,4; 8,0; 0) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,1; 0,0,0,2; 0,4,0; 8,0; 0) $\mathbb{L}_{1,0}\Delta_{2}\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,1; 0,0,0,2; 4,8,0; 16,0; 0) $\mathbb{L}_{6,1}$ (0,0,0,0,1; 0,0,1,0; 1,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{4,1}$ (0,0,0,0,1; 0,0,1,0; 2,0,0; 0,0; 0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{3}\Delta$ (0,0,0,0,1; 0,0,2,0; 0,4,0; 8,0; 0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{4,0}$ (0,0,0,0,1; 0,0,2,0; 2,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta^{2}\times\mathbb{L}_{1,1}\Delta$ (0,0,0,0,1; 0,0,2,0; 4,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,1; 0,0,2,0; 4,0,8; 0,16; 0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,1; 0,0,2,0; 4,8,0; 16,0; 0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{2}\Delta_{2}$ (0,0,0,0,1; 0,2,4,0; 8,16,0; 32,0; 0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}\Delta_{3}$ (0,0,0,0,1; 2,4,8,0; 16,32,0; 64,0; 0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{3}\Delta$ (0,0,0,1,2; 0,0,0,2; 0,0,4; 0,8; 0) $\mathbb{L}_{1,1}^{2}\Delta^{2}$ (0,0,0,1,2; 0,0,0,2; 0,0,4; 8,0; 0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{4,1}$ (0,0,0,1,2; 0,0,0,2; 1,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta^{2}$ (0,0,0,1,2; 0,0,0,2; 4,0,0; 0,0; 0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,1,2; 0,0,0,2; 4,0,8; 0,16; 0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,0,1,2; 0,0,0,4; 8,0,0; 0,0; 16) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{2}\Delta_{2}$ (0,0,0,1,2; 0,0,4,8; 16,0,0; 0,0; 32) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,1,2; 0,4,0,0; 8,0,0; 0,0; 16) $\mathbb{L}_{1,0}\Delta_{2}\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,1,2; 4,8,0,0; 16,0,0; 0,0; 32) $\mathbb{L}_{1,1}^{2}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,1,2,4; 0,0,0,4; 0,0,8; 16,0; 0) $\mathbb{L}_{1,1}\Delta^{3}$ (0,0,1,2,4; 0,0,0,4; 0,2,0; 0,0; 0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,1,2,4; 0,0,0,4; 0,8,0; 16,0; 0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{2}\Delta_{2}$ (0,0,1,2,4; 0,0,0,4; 8,16,0; 32,0; 0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta_{3}$ (0,0,1,2,4; 8,0,0,0; 0, 0,0; 16, 32; 64) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}\Delta_{3}$ (0,1,2,4,8; 0,0,0,8; 16,32,0; 64,0; 0) $\mathbb{L}_{1,1}\Delta_{2}^{2}$ (0,1,2,4,8; 0,0,4,8; 16,0,0; 0,0; 32)
normal quaternionic matrices of order 6, level >2>2
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{3}\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,0; 0,0,0,1; 0,2,0; 4,0; 0) $\mathbb{L}_{3}\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,0; 0,0,0,1; 2,4,0; 8,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,0; 0,0,0,2; 0,0,4; 8,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,0; 0,0,0,2; 4,8,0; 16,0; 0) $\mathbb{L}_{3}\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,2; 0,0,0,2; 0,0,4; 1,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,2; 0,0,0,2; 0,0,4; 8,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,2; 0,0,0,2; 0,4,0; 8,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,2; 0,0,0,2; 4,8,0; 16,0; 0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1})\Delta$ (0,0,0,0,2; 0,0,0,4; 0,0,8; 16,0; 0) $\mathbb{L}_{1,0}^{2}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (0,0,0,0,2; 0,0,0,4; 0,8,0; 16,0; 0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta\times\mathbb{L}_{1,0}\Delta^{2}$ (0,0,0,0,2; 0,0,0,4; 8,16,0; 32,0; 0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{3}\Delta$ (0,0,0,0,2; 0,0,1,4; 8,0,0; 0,0; 16) $\mathbb{L}_{1}\times\mathbb{L}_{5}$ (0,0,0,0,2; 0,0,2,0; 2,0,0; 0,0; 0) $\mathbb{L}_{3}^{2}$ (0,0,0,0,2; 0,0,2,0; 3,0,0; 0,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{3}\times\mathbb{L}_{1,0}\Delta$ (0,0,0,0,2; 0,0,2,0; 4,0,0; 0,0; 0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{4,0}$ (0,0,0,0,2; 0,0,4,0; 4,0,0; 0,0; 0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,0}\Delta^{2}$ (0,0,0,0,2; 0,0,4,0; 8,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,2; 0,0,4,0; 8,0,16; 0,32; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,2,4; 0,0,0,4; 0,0,8; 2,0; 0) $\mathbb{L}_{1,1}^{2}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (0,0,0,2,4; 0,0,0,4; 0,0,8; 16,0; 0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1})\Delta$ (0,0,0,2,4; 0,0,0,4; 0,8,0; 16,0; 0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{3}\Delta$ (0,0,0,2,4; 0,0,0,4; 1,8,0; 16,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{3}$ (0,0,0,2,4; 0,0,0,4; 2,0,0; 0,0; 0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{4,1}$ (0,0,0,2,4; 0,0,0,4; 4,0,0; 0,0; 0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta$ (0,0,0,2,4; 0,0,0,4; 8,0,0; 0,0; 0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,2,4; 0,0,0,4; 8,16,0; 32,0; 0) $\mathbb{L}_{3}\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (0,0,0,2,4; 0,0,0,8; 2,0,0; 0,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (0,0,0,2,4; 0,0,0,8; 16,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2})\Delta$ (0,0,0,2,4; 0,0,0,8; 16,0,0; 0,0; 32) $\mathbb{L}_{3}\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,2,4; 0,0,2,4; 1,0,0; 0,0; 8) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,2,4; 0,0,2,4; 8,0,0; 0,0; 16) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,0,2,4; 0,0,2,8; 16,0,0; 0,0; 32) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{2}$ (0,0,0,2,4; 0,0,8,16; 32,0,0; 0,0; 64) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,1}\Delta^{2}$ (0,0,2,4,8; 0,0,0,8; 0,4,0; 0,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (0,0,2,4,8; 0,0,0,8; 0,16,0; 0,0; 0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2})\Delta$ (0,0,2,4,8; 0,0,0,8; 0,16,0; 32,0; 0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,2,4,8; 0,0,0,8; 2,4,0; 16,0; 0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,2,4,8; 0,0,0,8; 2,16,0; 32,0; 0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{2}$ (0,0,2,4,8; 0,0,0,8; 16,32,0; 64,0; 0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta^{2}$ (0,0,2,4,8; 0,0,0,16; 0,32,0; 0,8; 0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,2,4,8; 0,0,0,16; 2,4,0; 32,0; 0) $\mathbb{L}_{1}^{3}\times\mathbb{L}_{3}$ (0,0,2,4,8; 2,0,0,0; 0,0,0; 0,0; 0) $\mathbb{L}_{3}\times\mathbb{L}_{1}^{2}\Delta$ (0,0,2,4,8; 2,0,0,0; 0,0,0; 0,0; 16)